Two forces of magnitudes 6.0N 4.0N act on a 3.2kg body. What is the acceleration produced when these forces are acting in the same direction? What is the acceleration produced when these forces are opposite directed
1
Expert's answer
2017-06-10T05:45:11-0400
When forces are acting in the same direction the acceleration
a=(F_1+F_2)/m=(6.0+4.0)/3.2=3.125 m/s^2 .
When forces are opposite directed
a=(F_1-F_2)/m=(6.0-4.0)/3.2=0.625 m/s^2 .

Answers:
(1) 3.125 m/s^2 ;
(2) 0.625 m/s^2 .